OpenStudy (anonymous):

Multiply: (1-sin x)^2 How do i do this? Multiply: (1-sin x)^2 How do i do this? @Mathematics

myininaya (myininaya):

\[(1-\sin(x))^2=1-2\sin(x)+\sin^2(x)\] recall \[(a-b)^2=a^2-2ab+b^2\]